HIGH RESOLUTION OVERTONE SPECTROSCOPY OF ACETYLENE VAN DER WAALS COMPLEXES

\underline{K. DIDRICHE}, C. LAUZIN, T. FOLDES, X. DE GHELLINCK, M. HERMAN, Service de Chimie quantique et Photophysique CP160/09, Faculté des Sciences, Université Libre de Bruxelles (U.L.B.), Av. Roosevelt, 50, B-1050, Bruxelles, Belgium.

The latest upgrades and results from the FANTASIO+ set-up^{a} (for ’Fourier trANsform, Tunable, diode and quadrupole mAss spectrometers interfaced to a Supersonic expansIOn’) are presented. Thanks to the improvements of the calibration (better than ±1 × 10^{−3} cm^{−1} accuracy) and the ring-down time (130 μs), the spectrum of C_{2}H_{2}-N_{2} around 6500 cm^{−1} was recorded as well as absorption bands due to combined excitations of intra and intermolecular modes of C_{2}H_{2}-Ar, C_{2}H_{2}-CO_{2} and C_{2}H_{2}-N_{2}O. A hot absorption band of the van der Waals complex C_{2}H_{2}-Ar was also observed. The first analyses of these spectra will be presented.

Along the development of the FANTASIO+ set-up, a new injection system to probe samples liquid at STP conditions, and a pulse injector are currently under implementation. Those upgrades will allow the study of organic and prebiotic molecules and a better cooling of the gas in the expansion, respectively. The first results will also be presented.

aDidriche et al., Mol. Phys. 2010, 108, 2158-2164